# ThinkGeo [Uncaptioned image] : Evaluating Tool-Augmented Agents for Remote Sensing Tasks

### v1

@@ Abstract
Recent progress in large language models (LLMs) has enabled tool-augmented agents capable of solving complex real-world tasks through step-by-step reasoning. However, existing evaluations often focus on general-purpose or multimodal scenarios, leaving a gap in domain-specific benchmarks that assess tool-use capabilities in complex remote sensing use cases. We present ThinkGeo, an agentic benchmark designed to evaluate LLM-driven agents on remote sensing tasks via structured tool use and multi-step planning. Inspired by tool-interaction paradigms, ThinkGeo includes human-curated queries spanning a wide range of real-world applications such as urban planning, disaster assessment and change analysis, environmental monitoring, transportation analysis, aviation monitoring, recreational infrastructure, and industrial site analysis. Each query is grounded in satellite or aerial imagery and requires agents to reason through a diverse toolset. We implement a ReAct-style interaction loop and evaluate both open and closed-source LLMs (e.g., GPT-4o, Qwen2.5) on 436 structured agentic tasks. The benchmark reports both step-wise execution metrics and final answer correctness. Our analysis reveals notable disparities in tool accuracy and planning consistency across models. ThinkGeo provides the first extensive testbed for evaluating how tool-enabled LLMs handle spatial reasoning in remote sensing. Our code and dataset are publicly available.
@@ 1 Introduction
Recent advances in LLMs have enabled the emergence of tool-augmented agents, systems that can break down complex tasks into step-by-step plans, invoke external tools (e.g., vision modules, calculators, and code interpreters), and reason across intermediate states [35, 26]. This paradigm, popularized via ReAct-style frameworks [35], has shown promise in general-purpose settings through benchmarks like ToolBench [24], GAIA [21], and GTA [28], which evaluate agents on procedural correctness, tool use, and final task outcomes. However, these benchmarks largely focus on synthetic, open-domain, or web-grounded scenarios, leaving the question of agentic capability in precision-critical, spatially grounded domains, like remote sensing, largely unexplored.
Remote sensing (RS) has become critical to a wide range of applications, including environmental monitoring, urban infrastructure analysis, transportation analysis, disaster response, and land-use mapping, with an ever-growing stream of high-resolution imagery from earth observation (EO) satellites and drones [16]. While modern visual models for detection, segmentation, and change analysis have advanced significantly, current processing pipelines remain brittle and manually engineered across tasks. Integrating these capabilities into LLM-driven agents introduces unique challenges: it requires reasoning over geodetic metadata, spatial resolutions, temporal dynamics, and unit-aware calculations. Existing agentic benchmarks (e.g., GTA [28], GAIA [21]) do not address these demands; they are built around general-purpose or web-grounded images and queries, and lack the spatial fidelity and grounding required for geospatial workflows. As a result, there is a pressing need for a benchmark that evaluates how well tool-augmented agents can operate in remote sensing contexts, reasoning over real EO imagery, coordination of general-purpose visual tools, and handling spatially grounded multi-step tasks.
@@
  • Task Suite & Dataset: A curated set of 436 agentic tasks paired with 311 medium to high-resolution optical EO images, spanning urban, environmental, transportation, aviation, industrial, change detection, and disaster-related scenarios. Examples are shown in Figure 1.
@@
API-Bench [23]
ToolBench [24]
GAIA [21]
APIBank [19]
m&m’s [20]
GTA [28]
@@
2 Related Work
Tool-augmented LLM Agents and Benchmarks: Integrating large language models (LLMs) with executable tools has recently become a central focus in agent research. Early work presented tool use as an alternating planning and execution. ReAct, for instance, interleaves "thought" tokens with structured tool calls, enabling a single LLM to both reason and act [35]. Subsequent systems generalized this idea to larger tool repertoires. HuggingGPT employs a GPT controller to select from hundreds of vision, speech, and language models exposed as functions [26], while Visual ChatGPT and MM-ReAct demonstrate analogous pipelines for multimodal perception tasks [30, 34]. To measure tool-use proficiency, several benchmarks have been proposed. ToolBench, APIBench, and API-Bank evaluate single-step API invocation within synthetic prompts [24, 23, 19]; m&m’s extends this to multi-step multimodal settings [20]. More recently, GAIA [21] and GTA [28] introduced human-written, step-implicit tasks paired with executable tool chains, revealing substantial performance gaps: GPT-4 completes fewer than half of GTA queries once real tools and intermediate checks are enforced. MLGym casts the agent problem into a Gym environment for open-ended AI-research workflows, highlighting long-horizon planning and code execution and also without geospatial imagery [22].
@@ 2 Related Work
Remote Sensing Agents: Recent efforts to extend LLM agents into EO have produced diverse tool-augmented pipelines, yet planning transparency, and step-level reasoning fidelity remain limited. Remote Sensing ChatGPT [9] and RS-Agent [33] represent early vision-language pipelines that chain pretrained detectors, segmenters, and geospatial utilities under GPT-based planners. However, they typically report only final answer accuracy, omitting structured ReAct-style trace evaluation or step-wise error attribution. TreeGPT and GeoMap-Agent [7, 11] introduce domain-specific agents for forestry and geological mapping, respectively. While these systems operate over visual maps and structured visual inputs, they rely on template-grounded or qualitative responses and do not implement formal multi-step evaluation. UnivEARTH [16], by contrast, employs a purely language-based framework that requires LLMs to generate valid Google Earth Engine (GEE) code, revealing that over 58% of completions fail to execute and that even the best agents answer only around 33% of geospatial queries correctly. Together, these works suggest that while EO agents can interface with rich toolsets, failures in tool selection, argument grounding, and spatial unit reasoning persist, underscoring the need for benchmarks that explicitly evaluate tool-level correctness alongside geospatial task outcomes for geospatial tasks.
@@
Evaluation Protocols: Early benchmarks for tool-augmented LLMs, such as ToolBench [24], APIBench [23], and API-Bank [19], primarily evaluate single-step tool usage in synthetic or isolated API call settings. While useful for measuring basic tool and argument prediction, these setups lack support for multi-tool planning, intermediate tracking, or long-horizon reasoning. To address these limitations, GTA [28] presents a tightly scoped yet richly instrumented benchmark requiring sequential tool usage across perception, logic, operation, and generation modules. GTA adopts a ReAct-style interface and introduces fine-grained supervision for each agent step, reporting metrics like ToolAcc, ArgAcc, StepAcc, and final answer correctness, thereby uncovering latent failure modes in tool selection and planning. Complementing this, MLGym [22] reframes agent evaluation as multi-task episodic learning within a Gym-style environment, simulating end-to-end ML workflows (e.g., training, evaluation, reporting) that demand persistent memory and adaptive behavior.
@@ 3 ThinkGeo Benchmark
3.1 Design Goals
Geospatial Reasoning: Tasks are modeled after real-world use cases in environmental monitoring and disaster response. Queries reflect practical challenges such as measuring metre-scale distances between objects, counting structures within spatial buffers, and identifying attributes of flooded or damaged buildings. These tasks require fine spatial fidelity, unit-based reasoning, and visual attribute grounding, capabilities often overlooked in existing benchmarks such as GAIA [21] or ToolBench [24].
@@ 3.1 Design Goals
Step-Implicit Tool Use: Unlike benchmarks where tool use is predefined or explicitly mentioned (e.g., APIBench [23]), ThinkGeo presents step-implicit, tool-implicit queries. Prompts do not reference tools by name; agents must infer which modules (e.g., perception, logic, operation) are needed and in what order. This design promotes true agentic planning and aligns with ReAct-style decision traces as used in GTA [28].
@@ 3.2 Use Case Categories
ThinkGeo is organized into seven primary categories, each reflecting critical application domains within the remote sensing ecosystem. These include Urban Planning, Disaster Assessment & Change Analysis, Environmental Monitoring, Transportation Analysis, Aviation Monitoring, Recreational Infrastructure, and Industrial Sites. Fig. 2(a) shows the use case taxanomy.
Each category encapsulates a range of spatially grounded, tool-invoking subtasks inspired by operational workflows in urban analytics, environmental science, and infrastructure planning:
@@ 3.3 Query Construction Pipeline
Step 1: Data Sampling & Guidelines. We begin by curating high-quality remote sensing samples from diverse RS datasets. Domain experts provide task-specific guidelines and generate initial reference queries to bootstrap the selection process. Step 2: Authoring ReAct Format. Using these guidelines, annotators manually inspect RS imagery, identify relevant objects and spatial relationships, and construct natural language queries following the ReAct format [35]. This involves composing a user query that implies multiple reasoning steps, manually annotating missing elements where needed, and generating a semi-structured dialog trace (thoughts, tool calls, observations, answers). The process is supported by a script built on the OpenAI GPT API. This script leverages per-image metadata (e.g., object types, GSD, bounding boxes) and tool definitions to generate diverse, tool-requiring prompts. Used prompts in the query construction pipeline are given in the supplementary material. Step 3: Validation. All generated samples are verified through a two-stage validation protocol. First, expert reviewers assess the semantic correctness, relevance, and alignment with the toolset. Second, we apply script-based checks to validate tool argument consistency, dialog structure, and completeness. Invalid samples are manually refined and corrected before inclusion in the final dataset. Fig. 3 shows the end-to-end dataflow for constructing our ThinkGeo benchmark.
@@
  • Difficulty Annotation: We categorized the queries into easy and hard levels based on two criteria: the number of complex keywords present and the number of reasoning steps required. Queries containing terms such as "estimate," "compare," "distribution," "count," "area," "how many," "orientation," and "proximity," along with more reasoning steps, were considered harder. To organize the queries, we sorted them based on the count of complex keywords and the number of steps. Queries appearing earlier in the sorted list, with fewer complex keywords and shorter reasoning steps, were labeled as easy, while the rest were classified as hard. After difficulty annotations, we have 281 easy queries while 155 hard queries. This sorting strategy provides a simple yet effective way to separate queries by their semantic and procedural complexity.
  • Inference-Aligned Prompting: Prompts are designed such that the agent must recover the reasoning chain without relying on field names or explicit tool indicators, promoting alignment with real-world, instruction-following behavior during inference. This adheres to the tool-implicit design philosophy established in agentic benchmarks like GTA [28].
@@ 3.4 Source RS Datasets
To construct the ThinkGeo benchmark, we leverage a diverse set of publicly available remote sensing datasets (Table 2) spanning various domains: DOTA [31], NWPU-VHR-10 [5], UCAS-AOD [36], and iSAID [29] support transportation and aviation-related tasks; FloodNet [25] and xBD [10] contribute flood-specific and temporal disaster imagery; AID [32] covers urban and industrial scenes; and the Global Dumpsite Dataset [27] addresses environmental monitoring. Original images are reused, while task-specific annotations are added where the datasets lack required labels.
Table 2: Remote sensing datasets used as image sources in the construction of the ThinkGeo benchmark. These datasets span a wide range of applications, sensor resolutions, and annotation types. Notably, the agentic tasks defined on these images are newly annotated bottomup.
@@
DOTA [31] Monitoring Transport, Aviation, Infrastructure GSD, B-Box, Category (0.1–1)m/px 2021
NWPU-VHR-10 [5] Monitoring Transport, Aviation, Infrastructure B-Box, Category (0.5–2)m/px 2023
UCAS-AOD [36] Monitoring Transport, Aviation B-Box, Category (0.5–2)m/px 2015
AID [32] Urban Planning, Monitoring Transport, Industr. Sites B-Box, Category (0.2–2)m/px 2017
iSAID [29] Monitoring Transport GSD, B-Box, Seg. Map, Pixel Count (0.1–1)m/px 2019
xBD [10] Disaster Assessment & Change Analysis GSD, B-Box, Category, Pixel Count (1–3.5)m/px 2019
FloodNet [25] Urban Planning, Disaster, Transport Analysis GSD, B-Box, Category, Seg. Map, Pix. Count (0.015–0.02)m/px 2020
Global-Dumpsite [27] Environmental Monitoring B-Box, Category (0.3-0.8)m/px 2023

### v2

@@ Abstract
Recent progress in large language models (LLMs) has enabled tool-augmented agents capable of solving complex real-world tasks through step-by-step reasoning. However, existing evaluations often focus on general-purpose or multimodal scenarios, leaving a gap in domain-specific benchmarks that assess tool-use capabilities in complex remote sensing use cases. We present ThinkGeo, an agentic benchmark designed to evaluate LLM-driven agents on remote sensing tasks via structured tool use and multi-step planning. Inspired by tool-interaction paradigms, ThinkGeo includes human-curated queries spanning a wide range of real-world applications such as urban planning, disaster assessment and change analysis, environmental monitoring, transportation analysis, aviation monitoring, recreational infrastructure, and industrial site analysis. Queries are grounded in satellite or aerial imagery, including both optical RGB and SAR data, and require agents to reason through a diverse toolset. We implement a ReAct-style interaction loop and evaluate both open and closed-source LLMs (e.g., GPT-4o, Qwen2.5) on 486 structured agentic tasks with 1,773 expert-verified reasoning steps. The benchmark reports both step-wise execution metrics and final answer correctness. Our analysis reveals notable disparities in tool accuracy and planning consistency across models. ThinkGeo provides the first extensive testbed for evaluating how tool-enabled LLMs handle spatial reasoning in remote sensing.
@@ 1 Introduction
Recent advances in LLMs have enabled the emergence of tool-augmented agents, systems that can break down complex tasks into step-by-step plans, invoke external tools (e.g., vision modules, calculators, and code interpreters), and reason across intermediate states Yao et al. (2023); Shen et al. (2023). This paradigm, popularized via ReAct-style frameworks Yao et al. (2023), has shown promise in general-purpose settings through benchmarks like ToolBench Qin et al. (2023), GAIA Mialon et al. (2023), and GTA Wang et al. (2024), which evaluate agents on procedural correctness, tool use, and final task outcomes. However, these benchmarks largely focus on synthetic, open-domain, or web-grounded scenarios, leaving the question of agentic capability in precision-critical, spatially grounded domains, like remote sensing, largely unexplored.
Remote sensing (RS) is critical to a wide range of applications, including environmental monitoring, urban infrastructure and transportation analysis, disaster response, and land-use mapping, with an ever-growing stream of high-resolution imagery from earth observation (EO) satellites and drones Kao et al. (2025). Despite advances in visual models for detection, segmentation, and change analysis, current processing pipelines remain brittle and manually engineered across tasks. Integrating these capabilities into LLM-driven agents demands reasoning over geodetic metadata, spatial resolutions, temporal dynamics, and unit-aware calculations. Existing agentic benchmarks (e.g., GTA Wang et al. (2024), GAIA Mialon et al. (2023)) do not address these demands; they are built around general-purpose or web-grounded images, lacking the spatial fidelity and grounding required for geospatial workflows. Consequently, there is a pressing need for a benchmark that evaluates tool-augmented agents in remote sensing contexts, for reasoning over real EO imagery, coordination of general-purpose visual tools, and handling spatially grounded multi-step tasks.
@@
  • Task Suite & Dataset: A curated set of 486 agentic tasks with 1,773 expert-verified reasoning steps over medium to high-resolution optical RGB (with 436 tasks) and SAR (with 50 tasks) images, spanning urban, environmental, transportation, aviation, industrial, change detection, and disaster-related scenarios. Examples are shown in Figure 1.
@@
API-Bench Patil et al. (2023)
ToolBench Qin et al. (2023)
GAIA Mialon et al. (2023)
APIBank Li et al. (2023)
m&m’s Ma et al. (2024)
GTA Wang et al. (2024)
@@
2 Related Work
Tool-augmented LLM Agents and Benchmarks: Integrating large language models (LLMs) with executable tools has recently become a central focus in agent research. Early work presented tool use as an alternating planning and execution. ReAct, for instance, interleaves ”thought” tokens with structured tool calls, enabling a single LLM to both reason and act Yao et al. (2023). Subsequent systems generalized this idea to larger tool repertoires. HuggingGPT employs a GPT controller to select from hundreds of vision, speech, and language models exposed as functions Shen et al. (2023), while Visual ChatGPT and MM-ReAct demonstrate analogous pipelines for multimodal perception tasks Wu et al. (2023); Yang et al. (2024). To measure tool‐use proficiency, several benchmarks have been proposed. ToolBench, APIBench, and API-Bank evaluate single-step API invocation within synthetic prompts Qin et al. (2023); Patil et al. (2023); Li et al. (2023); m&m’s extends this to multi-step multimodal settings Ma et al. (2024). More recently, GAIA Mialon et al. (2023) and GTA Wang et al. (2024) introduced human-written, step-implicit tasks paired with executable tool chains, revealing substantial performance gaps: GPT-4 completes fewer than half of GTA queries once real tools and intermediate checks are enforced. MLGym casts the agent problem into a Gym environment for open-ended AI-research workflows, highlighting long-horizon planning and code execution and also without geospatial imagery Nathani et al. (2025).
@@ 2 Related Work
Remote Sensing Agents: Recent efforts to extend LLM agents into EO have produced diverse tool-augmented pipelines, yet planning transparency, and step-level reasoning fidelity remain limited. Remote Sensing ChatGPT Guo et al. (2024) and RS-Agent Xu et al. (2024) represent early vision-language pipelines that chain pretrained detectors, segmenters, and geospatial utilities under GPT-based planners. However, they typically report only final answer accuracy, omitting structured ReAct-style trace evaluation or step-wise error attribution. TreeGPT and GeoMap-Agent Du et al. (2023); Huang et al. (2024) introduce domain-specific agents for forestry and geological mapping, respectively. While these systems operate over visual maps and structured visual inputs, they rely on template-grounded or qualitative responses and do not implement formal multi-step evaluation. UnivEARTH Kao et al. (2025), by contrast, employs a purely language-based framework that requires LLMs to generate valid Google Earth Engine (GEE) code, revealing that over 58% of completions fail to execute and that even the best agents answer only around 33% of geospatial queries correctly. Together, these works suggest that while EO agents can interface with rich toolsets, failures in tool selection, argument grounding, and spatial unit reasoning persist, underscoring the need for benchmarks that explicitly evaluate tool-level correctness alongside geospatial task outcomes.
@@
Evaluation Protocols: Early benchmarks for tool-augmented LLMs, such as ToolBench Qin et al. (2023), APIBench Patil et al. (2023), and API-Bank Li et al. (2023), primarily evaluate single-step tool usage in synthetic or isolated API call settings. While useful for measuring basic tool and argument prediction, these setups lack support for multi-tool planning, intermediate tracking, or long-horizon reasoning. To address these limitations, GTA Wang et al. (2024) presents a tightly scoped yet richly instrumented benchmark requiring sequential tool usage across perception, logic, operation, and generation modules. GTA adopts a ReAct-style interface and introduces fine-grained supervision for each agent step, reporting metrics like ToolAcc, ArgAcc, StepAcc, and final answer correctness, thereby uncovering latent failure modes in tool selection and planning. Complementing this, MLGym Nathani et al. (2025) reframes agent evaluation as multi-task episodic learning within a Gym-style environment, simulating end-to-end ML workflows (e.g., training, evaluation, reporting) that demand persistent memory and adaptive behavior.
@@ 3 ThinkGeo Benchmark
3.1 Design Goals
Geospatial Reasoning: Tasks are modeled after real-world use cases in environmental monitoring and disaster response. Queries reflect practical challenges such as measuring metre-scale distances, counting structures within spatial buffers, and identifying features of damaged buildings. These tasks require fine spatial fidelity, unit-based reasoning, and visual attribute grounding, capabilities often overlooked in existing benchmarks such as GAIA Mialon et al. (2023) or ToolBench Qin et al. (2023).
Step-Implicit Tool Use: Unlike benchmarks where tool use is predefined or explicitly mentioned (e.g., APIBench Patil et al. (2023)), ThinkGeo presents step-implicit, tool-implicit queries. Prompts do not reference tools by name; agents must infer which modules (e.g., perception, logic, operation) are needed and in what order. This design promotes true agentic planning and aligns with ReAct-style decision traces as used in GTA Wang et al. (2024).
@@ 3.2 Use Case Categories
ThinkGeo is organized into seven primary categories, each reflecting critical application domains within the remote sensing ecosystem. These include Urban Planning, Disaster Assessment & Change Analysis, Environmental Monitoring, Transportation Analysis, Aviation Monitoring, Recreational Infrastructure, and Industrial Sites. Fig. 2(a) shows the use case taxanomy. Each category encapsulates a range of spatially grounded, tool-invoking subtasks inspired by operational workflows in urban analytics, environmental science, and infrastructure planning:
@@ 3.3 Query Construction Pipeline
Step 1: Data Sampling & Guidelines. We begin by curating high-quality samples from diverse RS datasets. Domain experts provide task-specific guidelines and generate initial reference queries to bootstrap the selection process. Step 2: Authoring ReAct Format. Using these guidelines, annotators manually inspect imagery, identify key objects and spatial relationships, and construct natural language queries following the ReAct format Yao et al. (2023). This involves composing a user query that implies multi-step reasoning, manually annotating missing elements, and generating a semi-structured dialog trace (thoughts, tool calls, observations, answers).
The process is supported by a script built on the OpenAI GPT API. This script leverages per-image metadata (e.g., object types, GSD, bounding boxes) and tool definitions to generate diverse, tool-requiring prompts. Used prompts in the query construction pipeline are given in the supplementary material. Step 3: Validation. All generated samples are verified through a two-stage validation protocol. First, expert reviewers assess the semantic correctness, relevance, and alignment with the toolset. Second, we apply script-based checks to validate tool argument consistency, dialog structure, and completeness. Invalid samples are manually refined and corrected before inclusion in the final dataset. Fig. 3 shows the end-to-end dataflow for constructing our ThinkGeo benchmark.
@@
  • Difficulty Annotation: We categorized the queries into easy and hard levels based on two criteria: the number of complex keywords present and the number of reasoning steps required. Queries containing terms such as ”estimate,” ”compare,” ”distribution,” ”count,” ”area,” ”how many,” ”orientation,” and ”proximity,” along with more reasoning steps, were considered harder. To organize the queries, we sorted them based on the count of complex keywords and the number of steps. Queries appearing earlier in the sorted list, with fewer complex keywords and shorter reasoning steps, were labeled as easy, while the rest were classified as hard. This sorting strategy provides a simple yet effective way to separate queries by their semantic and procedural complexity.
  • Inference-Aligned Prompting: Prompts are designed such that the agent must recover the reasoning chain without relying on field names or explicit tool indicators, promoting alignment with real-world, instruction-following behavior. This adheres to the tool-implicit design philosophy established in agentic benchmarks like GTA Wang et al. (2024).
@@ 3.4 Source RS Datasets
To construct the ThinkGeo benchmark, we leverage a diverse set of publicly available remote sensing datasets (Table 2) spanning various domains: DOTA Xia et al. (2018), NWPU-VHR-10 Bian (2023), UCAS-AOD Zhu et al. (2015), and iSAID Waqas Zamir et al. (2019) support transportation and aviation-related tasks; FloodNet Rahnemoonfar et al. (2021) and xBD Gupta et al. (2019) contribute flood-specific and temporal disaster imagery; AID Xia et al. (2017) covers urban and industrial scenes; and the Global Dumpsite Dataset Sun & Yin (2023) addresses environmental monitoring. To expand transport and object-specific coverage, we additionally incorporate SSDD Zhang et al. (2021) for maritime ship monitoring, SADD Zhang et al. (2022) for aviation and aircraft monitoring, and SIVED Lin et al. (2023) for ground vehicle detection. Original images are reused, while task-specific annotations are added where the datasets lack required labels.
Table 2: Remote sensing datasets used as image sources in the construction of the ThinkGeo benchmark. These datasets span a wide range of applications, sensor resolutions, and annotation types. Notably, the agentic tasks defined on these images are newly annotated bottom-up.
@@
Optical RGB Datasets
DOTA Xia et al. (2018) Monitoring Transport, Aviation, Infrastructure GSD, B-Box, Category (0.1–1)m/px 2021
NWPU-VHR-10 Bian (2023) Monitoring Transport, Aviation, Infrastructure B-Box, Category (0.5–2)m/px 2023
UCAS-AOD Zhu et al. (2015) Monitoring Transport, Aviation B-Box, Category (0.5–2)m/px 2015
AID Xia et al. (2017) Urban Planning, Monitoring Transport, Industr. Sites B-Box, Category (0.2–2)m/px 2017
iSAID Waqas Zamir et al. (2019) Monitoring Transport GSD, B-Box, Seg. Map, Pixel Count (0.1–1)m/px 2019
xBD Gupta et al. (2019) Disaster Assessment & Change Analysis GSD, B-Box, Category, Pixel Count (1–3.5)m/px 2019
FloodNet Rahnemoonfar et al. (2021) Urban Planning, Disaster, Transport Analysis GSD, B-Box, Category, Seg. Map, Pix. Count (0.015–0.02)m/px 2020
Global-Dumpsite Sun & Yin (2023) Environmental Monitoring B-Box, Category (0.3-0.8)m/px 2023
SAR Datasets
SSDD Zhang et al. (2021) Monitoring Transport B-Box, Category (1–15)m/px 2021
SADD Zhang et al. (2022) Aviation B-Box, Category (0.5–3)m/px 2022
SIVED Lin et al. (2023) Monitoring Transport B-Box, Category (0.1-0.3)m/px 2023